Find the slope and y- intercept of the line -8x + 5y=15
Answer by stanbon(75887)

Find the slope and y- intercept of the line -8x + 5y=15
----------------
y = (8/5)x + 3
----
slope = 8/5
y-int = 3
------------
Cheers,
Stan H.